OpenStudy (anonymous):

show that the function given by y=(e^(-9t)+9)^-1 satisfies y'+81y^2=9y and y(0)=1/10 Any help would be greatly appreciated!!

OpenStudy (blockcolder):

Find y' and substitute it into y'+81y^2 and do stuff to make it become 9y.

OpenStudy (anonymous):

so y' = -81y^2 ? then take the derivative?

OpenStudy (blockcolder):

What? No, first find y' since you have an equation for y then substitute in y'+81y^2, not equate it to 0.

OpenStudy (anonymous):

I think it's like: \[\LARGE y=\left( e^{-9t}+9\right)^{-1}\] and you have to find it's derivative... to show that it's the same to \[\LARGE y'+81y^2=9y\] or\[\LARGE y'=9y-81y^2\] the to find y' in point x=0 to see if it's 1/10 ...

OpenStudy (anonymous):

using the chain rule? oh gosh... I think my brain is slightly fried! Sorry, I must be tired...

OpenStudy (anonymous):

but what you're going to do with "e" ... it doesn't just "vanish" , and you don't have it in the second equation.. ! \[\LARGE y'=-1\cdot \left( e^{-9t}+9\right)^{-2}\cdot (e^{-9t}+9)'\] ...\[\LARGE y'=- \left( e^{-9t}+9\right)^{-2}\cdot (e^{-9t})'\] \[\LARGE y'=-(e^{-9t}+9)^{-2}\cdot (-9e^{-9t})\] \[\LARGE y'=9e^{-9t}(e^{-9t}+9)^{-2}\] \[\LARGE y'=\frac{9}{ e^{9t}\cdot (e^{-9t}+9 )^2} \] and this is UGLY....

OpenStudy (anonymous):

I'm sure there's something else required to do...

OpenStudy (anonymous):

hmmm.... thats a lot further than I got... I will study what you have written and see where it takes me. Thanks so much!! :)

OpenStudy (anonymous):

... but I think there's something else we should do, then how does this fit to the second equation given... ? ... I'm lost!

OpenStudy (anonymous):

ha! You and me both! This assignment is due Wednesday and Im stuck on this one darn question! Im pulling my hair out over it....

OpenStudy (anonymous):

IF .. (I can't call it a solution or something) that what I've done above it's correct.. then substituting.. \[\LARGE f(t)=\frac{9}{ e^{9t}\cdot (e^{-9t}+9 )^2}\] \[\LARGE f(0)=\frac{9}{ e^{9\cdot 0}\cdot (e^{-9\cdot 0}+9 )^2}\] \[\LARGE f(0)=\frac{9}{ e^{0}\cdot (e^{0}+9 )^2}\] \[\LARGE f(0)=\frac{9}{ 1\cdot (1+9 )^2}\] \[\LARGE f(0)=\frac{9}{ 100 }\] ... :\

OpenStudy (anonymous):

that's not correct either...

OpenStudy (anonymous):

hmmn...not to worry. I will keep at it. You've given me a few ideas to try so thank you! :)

OpenStudy (anonymous):

hold on...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

should the first step be ...\[y'=-1(e ^{-9t}+9)^{-2}\times-9e ^{-9t}\]

OpenStudy (anonymous):

yeah you just passed through the "shortcut" \[\LARGE (e^u)'=e^u\cdot u' \] which is the same...

OpenStudy (anonymous):

yep, so where to next?

OpenStudy (anonymous):

wait for someone... I don't know if we actually have to do this !!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

is the next step...\[y'+81y^2=-(e ^{-9t}+9)^{-2}\times-9e ^{-9t}+9(e ^{-9t}+9)^{-2}\]

OpenStudy (blockcolder):

You're actually pretty close. \[\large y'=\frac{9}{e^{9t}(e^{-9t}+9)^2}\\ \large 81y^2=\frac{81}{(e^{-9t}+9)^{2}}\\ \large y'+81y^2=\frac{9+81e^{9t}}{e^{9t}(e^{-9t}+9)^2}=\frac{9e^{9t}(e^{-9t}+9)}{e^{9t}(e^{-9t}+9)^2}=\frac{9}{e^{-9t}+9}=9y\]

OpenStudy (anonymous):

Thanks!!!

OpenStudy (dumbcow):

interesting ok i went ahead and solved the differential equation and got a different y \[y = \frac{e^{9t}}{1+9e^{9t}}\]

OpenStudy (blockcolder):

One last thing: Make sure that y(0)=1/10. \[y(0)=(e^0+9)^{-1}=(1+9)^{-1}=\frac{1}{10}\] @dumbcow Divide numerator and denominator by \(e^{-9t}\) to get the other y.

OpenStudy (dumbcow):

ahh thanks...anyway if you know how to solve diff equ its easier i think than finding derivative and squaring and plugging everyhting in

OpenStudy (anonymous):

\[ y(t)=\frac{1}{e^{-9 t}+9}\\ y'(t) = \frac{9 e^{-9 t}}{\left(e^{-9 t}+9\right)^2}\\ y'(t) + 81 y(t)^2 = \frac{9 e^{-9 t}}{\left(e^{-9 t}+9\right)^2}+\frac{81}{\left(e^{-9 t}+9\right)^2}= \frac{9 \left(e^{-9 t}+9\right)}{\left(e^{-9 t}+9\right)^2}\\ \frac{9}{e^{-9 t}+9} =9 y(t)\\ y(0) = \frac 1 {10} \]